# 2]International Research Center for Intelligent Perception and Computation, Xi’an 710071, China

### v1

@@ Abstract
Real-world data are long-tailed, the lack of tail samples leads to a significant limitation in the generalization ability of the model. Although numerous approaches of class re-balancing perform well for moderate class imbalance problems, additional knowledge needs to be introduced to help the tail class recover the underlying true distribution when the observed distribution from a few tail samples does not represent its true distribution properly, thus allowing the model to learn valuable information outside the observed domain. In this work, we propose to leverage the geometric information of the feature distribution of the well-represented head class to guide the model to learn the underlying distribution of the tail class. Specifically, we first systematically define the geometry of the feature distribution and the similarity measures between the geometries, and discover four phenomena regarding the relationship between the geometries of different feature distributions. Then, based on four phenomena, feature uncertainty representation is proposed to perturb the tail features by utilizing the geometry of the head class feature distribution. It aims to make the perturbed features cover the underlying distribution of the tail class as much as possible, thus improving the model’s generalization performance in the test domain. Finally, we design a three-stage training scheme enabling feature uncertainty modeling to be successfully applied. Experiments on CIFAR-10/100-LT, ImageNet-LT, and iNaturalist2018 show that our proposed approach outperforms other similar methods on most metrics. In addition, the experimental phenomena we discovered are able to provide new perspectives and theoretical foundations for subsequent studies. The code will be available at https://github.com/mayanbiao1234/Geometric-metrics-for-perceptual-manifolds-in-deep-neural-networks/
@@
Long-Tailed Classification, Representational learning, Geometric prior knowledge
@@ 1 Introduction
Deep learning has made significant progress in image classification, image segmentation, and other fields benefiting from artificially annotated large-scale datasets. However, real-world data tends to follow a long-tailed distribution [38], and unbalanced classes introduce bias into machine learning models. Numerous approaches have been proposed to mitigate the model bias, such as class re-balancing [6, 13, 56, 24, 31, 32], information augmentation [4, 21, 50, 55, 19] and network structure design [21, 51, 56]. However, the above approach does not work effectively in all cases, and the generalization ability of the model will be greatly limited when the samples of the tail class do not accurately represent its true distribution. We discuss two cases of the relationship between the observed and true distributions of the tail classes [5].
@@
In case 1, although the sample size of the tail classes is small, these samples represent the true data distribution. The main reason for the degradation of the model performance is that at each sampling, samples from the tail class are used with a small probability to calculate loss and update parameters, resulting in inadequate learning of the tail class by the model. Faced with this situation, existing data augmentation methods [4, 47], undersampling [41], oversampling [42, 50], and rebalancing loss [20, 27, 31] can reasonably improve performance. Combining decoupled training with the above approach can improve the performance of the model even further [1, 51]. However, rebalancing strategies, including numerous meta-learning methods [31, 32, 11], do not increase the information outside the observed training domain. Some meta-learning methods [19] used for information augmentation also struggle to provide valuable information due to weak guidance from the meta-data [48]. Additionally, CReST [44], a self-training-based approach, attempts to label new samples for the tail classes, but the introduced new samples are already well-recognized by the long-tailed model, resulting in very limited new information being introduced. In summary, when certain classes are severely underrepresented (i.e., case 2), these methods have difficulty finding the right direction for adjusting the decision boundaries, so improvements sometimes worsen [5, 52]. Therefore, we pursue to mine knowledge from the well-represented head classes to help recover the true distribution of the tail classes.
@@
In case 2, if the underlying true distribution of the tail classes cannot be recovered, then the model always fails to learn the correct decision boundaries. Even if the model achieves high recognition accuracy in the training set, it still fails to have satisfactory generalization performance when faced with test samples outside the training domain. Therefore, if the direction for recovering the true distribution of tail classes, such as the direction indicated by the three arrows in Figure 1c, can be found, the generalization ability of the model on the tail classes will be significantly improved. It is necessary to explore additional knowledge to guide the tail class to recover the true distribution. Transfer learning for long-tail classification aims to introduce new information to facilitate the model’s learning on tail classes, which can be classified into two categories [52]: head-to-tail knowledge transfer and model pretraining. Due to the natural scarcity of tail class data, model pretraining is not effectively applicable in long-tail scenarios [48, 52]. Therefore, most of the current research focuses on head-to-tail knowledge transfer. However, the motivations for previous head-to-tail knowledge transfer were limited to qualitative analysis or speculation [1, 14, 30, 39]. They assumed that all background information from head classes exists in tail classes, which is an unsupported assumption lacking evidence. Moreover, these methods neglected the potential diversity of foreground information in tail classes, leading to the possibility of inadequate recovery of the true distribution of tail classes. Our experimental comparisons also demonstrate that knowledge transfer supported by evidence outperforms the aforementioned methods. In conclusion, recovering the underlying true distribution with limited samples remains a meaningful challenge.
@@
It has been shown that the model bias is caused by the classifier and the long-tailed data does not unbalance the feature representation learning [13, 56]. Also considering that the dimensionality of the feature space is smaller than that of the sample space, we focus on recovering the true distribution of tail classes in the feature space. In this work, our main contributions are summarized as follows.
@@ 2.1 Class Rebalancing
The extreme imbalance in the number of samples in the long-tail data prevents the classification model from learning the distribution of the tail classes adequately, which leads to poor performance of the model on the tail classes. Therefore, methods to rebalance the number of samples and the losses incurred per class (i.e., resampling and reweighting) are proposed. Resampling methods are divided into oversampling and undersampling [3, 8, 9, 13, 40, 53]. The idea of oversampling is to randomly sample the tail classes to equalize the number of samples and thus optimize the classification boundaries. The undersampling methods balance the number of samples by randomly removing samples from the head classes. For example, [41] finds that training with a balanced subset of a long-tailed dataset is instead better than using the full dataset. In addition, [13, 56] fine-tune the classifier via a resampling strategy in the second phase of decoupled training. [42] continuously adjusts the distribution of resampled samples and the weights of the two-loss terms during training to make the model perform better. [50] employs the model classification loss from an additional balanced validation set to adjust the sampling rate of different classes.
@@
The purpose of reweighting loss is intuitive, and it is proposed to balance the losses incurred by all classes, usually by applying a larger penalty to the tail classes on the objective function (or loss function) [7, 10, 35, 36, 43, 54, 26, 24]. [31] proposes to adjust the loss with the label frequencies to alleviate class bias. [20] not only assigns weights to the loss of each class but also assigns higher weights to hard samples. Recent studies have shown that the effect of reweighting losses by the inverse of the number of samples is modest [27, 28]. Some methods that produce more “smooth" weights for reweighting perform better [24], such as taking the square root of the number of samples as the weight [27]. [6] attributes the better performance of this smoother method to the existence of marginal effects. In addition, [1] proposes to learn the classifier with class-balanced loss by adjusting the weight decay and MaxNorm in the second stage. DSB [24] and DCR [26], for the first time, examined the factors influencing model bias from a geometric perspective and proposed a rebalancing approach.
@@
Although class rebalancing methods are simple to implement, their limitations have been increasingly recognized in recent research [52, 29]. Class rebalancing methods merely increase the weight of the tail class loss without introducing additional knowledge to assist the tail classes, which often leads to overfitting of the tail classes and significantly compromises the model’s generalization performance [29]. Another limitation is that class rebalancing methods often improve tail class performance at the expense of sacrificing head class performance, making it challenging to handle data scarcity issues [52, 48]. As a result, more and more research is focusing on information augmentation.
@@ 2.2 Stage-wise training
Decoupling [13] first proposes to decouple the learning process on long-tail data into feature learning and classifier learning, and it finds that re-learning the balanced classifier can significantly improve the model performance. Further, BBN [56] combines the two-stage learning into a two-branch model. The two branches of the model share parameters, with one branch learning using the original data and the other learning using the resampled data. [5] decomposes the features into class-generic features and class-specific features, and it expands the tail class data by combining class-generic features of the head class with class-specific features of the tail class. [55] finds that augmenting data with Mixup in the first stage benefits feature learning and does negligible damage to classifiers trained using decoupling. [51] also observes that long-tailed data does not affect feature learning, and it proposes an adaptive calibration function for improving the cross-entropy loss. [12] considers the effect of noisy samples on the tail class and adaptively assigns weights to the tail class samples by meta-learning in the second stage.
The two-stage training pushes the decision boundary away from the augmented tail class distribution, thereby improving the performance of the tail classes. However, this may lead to excessive bias in the decision boundary and affect the head classes [49]. Therefore, we propose a three-stage training strategy. The first two stages are indistinguishable from decoupled training, while in the third stage, we fix the classifier parameters and fine-tune the feature extractor to adapt it to the improved classification boundary.
@@ 2.3 Head-to-tail knowledge transfer
Head-to-tail knowledge transfer is more relevant to our work than other methods. [49] and [22] were first proposed in the face recognition field to transfer variance between classes to augment classes with fewer samples. [49, 25] assumes that the feature distributions of each class are multivariate Gaussian, and the feature distributions of the common and under-represented classes have the same variance, the variance of the head class is used to estimate the distribution of the tail class. [22] assumes that the intra-class angle distribution follows a Gaussian distribution, transfers the intra-class angle distribution of features to the tail class, and constructs a “feature cloud" for each feature to extend the distribution of the tail class. Both [49] and [22] assume that certain statistical properties of head and tail classes are the same (i.e., variance and intra-class angular distribution). However, these assumptions lack supporting evidence, which limits their performance in long-tail recognition.
Similar to the adversarial attack, [14] proposes to transform some of the head samples into tail samples through perturbation-based optimization to achieve tail class augmentation. However, this method exploits the vulnerabilities of deep neural networks to generate samples that mislead the model, and these samples do not exist as tail-class samples in reality. As a result, it cannot effectively help the tail classes recover their underlying distribution. [5] decomposes the features of each class into class-generic features and class-specific features. During training, the class-specific features of the tail class are fused with the generic features of the head class to generate new features to expand the tail class. This idea is similar to data augmentation in image space, such as Cutmix. [39] dynamically estimates a set of centers for each class, and then calculates the displacement between the head class feature and the corresponding nearest intra-class center. This displacement is used to combine with the tail class centers to generate new features, thereby increasing the feature diversity of the tail class. [21] proposes to transfer the geometric information of the feature distribution boundaries of the head class to the tail class by enhancing the weights of the tail class classifier. The recently proposed CMO [30] considers that the image of the head class has a rich background, so the image of the tail class can be pasted directly onto the background image of the head class to increase the richness of the tail class. Overall, the above-mentioned research assumes that all background information of head classes exists in tail classes, which is an unsupported assumption. It also overlooks the potential diversity of tail class foreground information, which may lead to ineffective recovery of the true distribution of tail classes.
@@ 3.1 The Geometry of Data Distribution
Definition 1 (The geometry of data distribution).
@@ 3.2 Similarity Measure of Geometry
Definition 2 (Similarity metric between geometry).
@@
$\begin{split}S(GD_{X_{1}},GD_{X_{2}})=\sum_{i=1}^{P}\langle\xi_{X_{1}}^{i},\xi_{X_{2}}^{i}\rangle=\sum_{i=1}^{P}{\xi_{X_{1}}^{i}}^{T}\xi_{X_{2}}^{i}.\end{split}$
@@
it is clear that the geometry of the two distributions in Figure 8 is more similar compared to the two distributions shown in Figure 9. The details are described in Appendix A.
@@ 3.3 Four Discoveries about the Geometry of the Feature Distribution
We investigated the relationship between class similarity and the geometry similarity of class distributions on two benchmark datasets: Fashion MNIST [45] and CIFAR-10 [15]. ResNet-18 [46] was adopted as the backbone network and various training schemes were applied to make the performance of ResNet-18 on the two datasets comparable to the state-of-the-art results (See Appendix B for details). First, the similarity between all classes on the two datasets is calculated and ranked. Then, we extracted $64$-dimensional features of all samples from both datasets using ResNet18 and calculated the geometry of all class feature distributions. Based on this, we summarize further experiments and findings as follows.
@@ 3.3.1 Phenomenon 1
As shown in Figure 2, features were extracted using trained ResNet-18 on MNIST [16], Fashion MNIST and CIFAR-10. We find the sum of the eigenvalues corresponding to the first five eigenvectors that are used to represent the geometry of the distribution can reach more than $80\%$ of the sum of all eigenvalues, which means that most of the information of the data distribution can be recovered along the first five eigenvectors.
Figure 2: The ratio of the sum of the top five eigenvalues to the sum of all eigenvalues after eigendecomposition for the feature embeddings of all classes in the three datasets. The horizontal coordinates are the indexes of the classes, and the specific class names are in Appendix D.

### v2

@@ Abstract
Real-world data are long-tailed, the lack of tail samples leads to a significant limitation in the generalization ability of the model. Although numerous approaches of class re-balancing perform well for moderate class imbalance problems, additional knowledge needs to be introduced to help the tail class recover the underlying true distribution when the observed distribution from a few tail samples does not represent its true distribution properly, thus allowing the model to learn valuable information outside the observed domain. In this work, we propose to leverage the geometric information of the feature distribution of the well-represented head class to guide the model to learn the underlying distribution of the tail class. Specifically, we first systematically define the geometry of the feature distribution and the similarity measures between the geometries, and discover four phenomena regarding the relationship between the geometries of different feature distributions. Then, based on four phenomena, feature uncertainty representation is proposed to perturb the tail features by utilizing the geometry of the head class feature distribution. It aims to make the perturbed features cover the underlying distribution of the tail class as much as possible, thus improving the model’s generalization performance in the test domain. Finally, we design a three-stage training scheme enabling feature uncertainty modeling to be successfully applied. Experiments on CIFAR-10/100-LT, ImageNet-LT, and iNaturalist2018 show that our proposed approach outperforms other similar methods on most metrics. In addition, the experimental phenomena we discovered are able to provide new perspectives and theoretical foundations for subsequent studies. The code will be available at https://github.com/mayanbiao1234/Geometric-metrics-for-perceptual-manifolds
@@
Long-Tailed Classification Representational learning Geometric prior knowledge
@@ 1 Introduction
Deep learning has made significant progress in image classification, image segmentation, and other fields benefiting from artificially annotated large-scale datasets. However, real-world data tends to follow a long-tailed distribution paper33 , and unbalanced classes introduce bias into machine learning models. Numerous approaches have been proposed to mitigate the model bias, such as class re-balancing paper6 ; paper12 ; paper50 ; paper53 ; paper27 ; learning , information augmentation paper4 ; paper20 ; paper44 ; paper49 ; paper18 and network structure design paper20 ; paper45 ; paper50 . However, the above approach does not work effectively in all cases, and the generalization ability of the model will be greatly limited when the samples of the tail class do not accurately represent its true distribution. We discuss two cases of the relationship between the observed and true distributions of the tail classes paper5 .
@@
In case 1, although the sample size of the tail classes is small, these samples represent the true data distribution. The main reason for the degradation of the model performance is that at each sampling, samples from the tail class are used with a small probability to calculate loss and update parameters, resulting in inadequate learning of the tail class by the model. Faced with this situation, existing data augmentation methods paper4 ; paper41 , undersampling paper36 , oversampling paper37 ; paper44 , and rebalancing loss paper19 ; paper23 ; paper27 can reasonably improve performance. Combining decoupled training with the above approach can improve the performance of the model even further paper1 ; paper45 . However, rebalancing strategies, including numerous meta-learning methods paper27 ; learning ; meta-rethinking , do not increase the information outside the observed training domain. Some meta-learning methods paper18 used for information augmentation also struggle to provide valuable information due to weak guidance from the meta-data paper42 . Additionally, CReST self-crest , a self-training-based approach, attempts to label new samples for the tail classes, but the introduced new samples are already well-recognized by the long-tailed model, resulting in very limited new information being introduced. In summary, when certain classes are severely underrepresented (i.e., case 2), these methods have difficulty finding the right direction for adjusting the decision boundaries, so improvements sometimes worsen paper5 ; paper46 . Therefore, we pursue to mine knowledge from the well-represented head classes to help recover the true distribution of the tail classes.
@@
In case 2, if the underlying true distribution of the tail classes cannot be recovered, then the model always fails to learn the correct decision boundaries. Even if the model achieves high recognition accuracy in the training set, it still fails to have satisfactory generalization performance when faced with test samples outside the training domain. Therefore, if the direction for recovering the true distribution of tail classes, such as the direction indicated by the three arrows in Figure 1c, can be found, the generalization ability of the model on the tail classes will be significantly improved. It is necessary to explore additional knowledge to guide the tail class to recover the true distribution. Transfer learning for long-tail classification aims to introduce new information to facilitate the model’s learning on tail classes, which can be classified into two categories paper46 : head-to-tail knowledge transfer and model pretraining. Due to the natural scarcity of tail class data, model pretraining is not effectively applicable in long-tail scenarios paper42 ; paper46 . Therefore, most of the current research focuses on head-to-tail knowledge transfer. However, the motivations for previous head-to-tail knowledge transfer were limited to qualitative analysis or speculation paper1 ; paper14 ; paper26 ; paper34 . They assumed that all background information from head classes exists in tail classes, which is an unsupported assumption lacking evidence. Moreover, these methods neglected the potential diversity of foreground information in tail classes, leading to the possibility of inadequate recovery of the true distribution of tail classes. Our experimental comparisons also demonstrate that knowledge transfer supported by evidence outperforms the aforementioned methods. In conclusion, recovering the underlying true distribution with limited samples remains a meaningful challenge.
@@
It has been shown that the model bias is caused by the classifier and the long-tailed data does not unbalance the feature representation learning paper12 ; paper50 . Also considering that the dimensionality of the feature space is smaller than that of the sample space, we focus on recovering the true distribution of tail classes in the feature space. In this work, our main contributions are summarized as follows.
@@ 2.1 Class Rebalancing
The extreme imbalance in the number of samples in the long-tail data prevents the classification model from learning the distribution of the tail classes adequately, which leads to poor performance of the model on the tail classes. Therefore, methods to rebalance the number of samples and the losses incurred per class (i.e., resampling and reweighting) are proposed. Resampling methods are divided into oversampling and undersampling paper3 ; paper8 ; paper9 ; paper12 ; paper35 ; paper47 . The idea of oversampling is to randomly sample the tail classes to equalize the number of samples and thus optimize the classification boundaries. The undersampling methods balance the number of samples by randomly removing samples from the head classes. For example, paper36 finds that training with a balanced subset of a long-tailed dataset is instead better than using the full dataset. In addition, paper12 ; paper50 fine-tune the classifier via a resampling strategy in the second phase of decoupled training. paper37 continuously adjusts the distribution of resampled samples and the weights of the two-loss terms during training to make the model perform better. paper44 employs the model classification loss from an additional balanced validation set to adjust the sampling rate of different classes.
@@
The purpose of reweighting loss is intuitive, and it is proposed to balance the losses incurred by all classes, usually by applying a larger penalty to the tail classes on the objective function (or loss function) paper7 ; paper10 ; paper30 ; paper31 ; paper38 ; paper48 ; paper52 ; paper53 . paper27 proposes to adjust the loss with the label frequencies to alleviate class bias. paper19 not only assigns weights to the loss of each class but also assigns higher weights to hard samples. Recent studies have shown that the effect of reweighting losses by the inverse of the number of samples is modest paper23 ; paper25 . Some methods that produce more “smooth” weights for reweighting perform better paper53 , such as taking the square root of the number of samples as the weight paper23 . paper6 attributes the better performance of this smoother method to the existence of marginal effects. In addition, paper1 proposes to learn the classifier with class-balanced loss by adjusting the weight decay and MaxNorm in the second stage. DSB paper53 and DCR paper52 , for the first time, examined the factors influencing model bias from a geometric perspective and proposed a rebalancing approach.
@@
Although class rebalancing methods are simple to implement, their limitations have been increasingly recognized in recent research paper46 ; paper54 . Class rebalancing methods merely increase the weight of the tail class loss without introducing additional knowledge to assist the tail classes, which often leads to overfitting of the tail classes and significantly compromises the model’s generalization performance paper54 . Another limitation is that class rebalancing methods often improve tail class performance at the expense of sacrificing head class performance, making it challenging to handle data scarcity issues paper46 ; paper42 . As a result, more and more research is focusing on information augmentation.
@@ 2.2 Stage-wise training
Decoupling paper12 first proposes to decouple the learning process on long-tail data into feature learning and classifier learning, and it finds that re-learning the balanced classifier can significantly improve the model performance. Further, BBN paper50 combines the two-stage learning into a two-branch model. The two branches of the model share parameters, with one branch learning using the original data and the other learning using the resampled data. paper5 decomposes the features into class-generic features and class-specific features, and it expands the tail class data by combining class-generic features of the head class with class-specific features of the tail class. paper49 finds that augmenting data with Mixup in the first stage benefits feature learning and does negligible damage to classifiers trained using decoupling. paper45 also observes that long-tailed data does not affect feature learning, and it proposes an adaptive calibration function for improving the cross-entropy loss. paper11 considers the effect of noisy samples on the tail class and adaptively assigns weights to the tail class samples by meta-learning in the second stage.
The two-stage training pushes the decision boundary away from the augmented tail class distribution, thereby improving the performance of the tail classes. However, this may lead to excessive bias in the decision boundary and affect the head classes paper43 . Therefore, we propose a three-stage training strategy. The first two stages are indistinguishable from decoupled training, while in the third stage, we fix the classifier parameters and fine-tune the feature extractor to adapt it to the improved classification boundary.
@@ 2.3 Head-to-tail knowledge transfer
Head-to-tail knowledge transfer is more relevant to our work than other methods. paper43 and paper21 were first proposed in the face recognition field to transfer variance between classes to augment classes with fewer samples. paper43 ; paper55 assumes that the feature distributions of each class are multivariate Gaussian, and the feature distributions of the common and under-represented classes have the same variance, the variance of the head class is used to estimate the distribution of the tail class. paper21 assumes that the intra-class angle distribution follows a Gaussian distribution, transfers the intra-class angle distribution of features to the tail class, and constructs a “feature cloud” for each feature to extend the distribution of the tail class. Both paper43 and paper21 assume that certain statistical properties of head and tail classes are the same (i.e., variance and intra-class angular distribution). However, these assumptions lack supporting evidence, which limits their performance in long-tail recognition.
Similar to the adversarial attack, paper14 proposes to transform some of the head samples into tail samples through perturbation-based optimization to achieve tail class augmentation. However, this method exploits the vulnerabilities of deep neural networks to generate samples that mislead the model, and these samples do not exist as tail-class samples in reality. As a result, it cannot effectively help the tail classes recover their underlying distribution. paper5 decomposes the features of each class into class-generic features and class-specific features. During training, the class-specific features of the tail class are fused with the generic features of the head class to generate new features to expand the tail class. This idea is similar to data augmentation in image space, such as Cutmix. paper34 dynamically estimates a set of centers for each class, and then calculates the displacement between the head class feature and the corresponding nearest intra-class center. This displacement is used to combine with the tail class centers to generate new features, thereby increasing the feature diversity of the tail class. paper20 proposes to transfer the geometric information of the feature distribution boundaries of the head class to the tail class by enhancing the weights of the tail class classifier. The recently proposed CMO paper26 considers that the image of the head class has a rich background, so the image of the tail class can be pasted directly onto the background image of the head class to increase the richness of the tail class. Overall, the above-mentioned research assumes that all background information of head classes exists in tail classes, which is an unsupported assumption. It also overlooks the potential diversity of tail class foreground information, which may lead to ineffective recovery of the true distribution of tail classes.
@@ 3.1 The Geometry of Data Distribution
Definition 1 (The geometry of data distribution)
@@ 3.2 Similarity Measure of Geometry
Definition 2 (Similarity metric between geometry)
@@
$\begin{split}S(GD_{X_{1}},GD_{X_{2}})=\sum_{i=1}^{P}\left|\langle\xi_{X_{1}}^{i},\xi_{X_{2}}^{i}\rangle\right|=\sum_{i=1}^{P}\left|{\xi_{X_{1}}^{i}}^{T}\xi_{X_{2}}^{i}\right|.\end{split}$
@@
it is clear that the geometry of the two distributions in Figure 8 is more similar compared to the two distributions shown in Figure 9. The details are described in Appendix Appendix A.
@@ 3.3 Four Discoveries about the Geometry of the Feature Distribution
We investigated the relationship between class similarity and the geometry similarity of class distributions on two benchmark datasets: Fashion MNIST paper39 and CIFAR-10 paper15 . ResNet-18 paper40 was adopted as the backbone network and various training schemes were applied to make the performance of ResNet-18 on the two datasets comparable to the state-of-the-art results (See Appendix Appendix B for details). First, the similarity between all classes on the two datasets is calculated and ranked. Then, we extracted $64$-dimensional features of all samples from both datasets using ResNet18 and calculated the geometry of all class feature distributions. Based on this, we summarize further experiments and findings as follows.
@@ 3.3.1 Phenomenon 1
As shown in Figure 2, features were extracted using trained ResNet-18 on MNIST paper16 , Fashion MNIST and CIFAR-10. We find the sum of the eigenvalues corresponding to the first five eigenvectors that are used to represent the geometry of the distribution can reach more than $80\%$ of the sum of all eigenvalues, which means that most of the information of the data distribution can be recovered along the first five eigenvectors.
Figure 2: The ratio of the sum of the top five eigenvalues to the sum of all eigenvalues after eigendecomposition for the feature embeddings of all classes in the three datasets. The horizontal coordinates are the indexes of the classes, and the specific class names are in Appendix Appendix D.